### v2

# Multimodal Medical Code Tokenizer

Abstract
Foundation models trained on patient electronic health records (EHRs) require tokenizing medical data into sequences of discrete vocabulary items. Existing tokenizers treat medical codes from EHRs as isolated textual tokens. However, each medical code is defined by its textual description, its position in ontological hierarchies, and its relationships to other codes, such as disease co-occurrences and drug-treatment associations. Medical vocabularies contain more than 600,000 codes with critical information for clinical reasoning. We introduce MedTok, a multimodal medical code tokenizer that uses the text descriptions and relational context of codes. MedTok processes text using a language model encoder and encodes the relational structure with a graph encoder. It then quantizes both modalities into a unified token space, preserving modality-specific and cross-modality information. We integrate MedTok into five EHR models and evaluate it on operational and clinical tasks across in-patient and out-patient datasets, including outcome prediction, diagnosis classification, drug recommendation, and risk stratification. Swapping standard EHR tokenizers with MedTok improves AUPRC across all EHR models, by 4.10% on MIMIC-III, 4.78% on MIMIC-IV, and 11.30% on EHRShot, with the largest gains in drug recommendation. Beyond EHR modeling, we demonstrate using MedTok tokenizer with medical QA systems. Our results demonstrate the potential of MedTok as a unified tokenizer for medical codes, improving tokenization for medical foundation models.
Medical Code, Tokenizer
1 Introduction
Figure 1: MedTok is a multimodal tokenizer for medical codes that combines text embeddings from code descriptions with graph-based representations of dependencies from ontologies and terminologies. It is a general tokenizer that sheds light on how optimized tokenization benefits transformer models in medicine.
Electronic health records (EHRs) are the backbone of modern healthcare, capturing a person’s health state with increasing precision across diverse modalities. Structured EHR data, encoded through standardized medical codes, support a wide range of applications, from personalized risk prediction (Goldstein et al., 2016; Yu et al., 2024b) and disease trajectory modeling (Jensen et al., 2017; Heumos et al., 2024) to emulation of clinical trials (Katsoulakis et al., 2024; Kraljevic et al., 2024). The cornerstone of structured EHRs is medical coding systems, which assign standardized alphanumeric codes to various aspects of patient health, including diseases, procedures, medications, and laboratory tests. These codes come from widely used terminologies such as ICD-9, ICD-10, SNOMED CT, CPT, and ATC, among others (Foley et al., 1992; Organization et al., 1988; Organization, 2004; Donnelly et al., 2006; Dotson, 2013; Miller & Britt, 1995). Although essential for interoperability, medical codes introduce challenges for models, particularly in the tokenization process, which transforms structured EHR data into token sequences that foundation models can process.
Transformer-based models for structured EHRs (Poulain & Beheshti, 2024; Yang et al., 2023b; Jiang et al., 2023b; Renc et al., 2024) rely on tokenizers to map raw data into discrete vocabulary items. However, standard tokenization strategies inherited from general-purpose language models fail to capture the complexity of medical codes, leading to six key challenges: (1) Scalability of medical vocabularies – Medical coding systems contain over 600,000 unique codes, far exceeding standard tokenizer capacities. Treating each code as a separate token leads to inefficient vocabulary expansion, increasing memory demands and fragmenting rare codes (e.g., splitting ”ICD9: 250.0” into arbitrary subwords). (2) Loss of hierarchical and relational structure – Many coding systems encode structured dependencies, such as ATC codes, which classify drugs based on pharmacological and chemical properties (Miller & Britt, 1995). Standard tokenizers, relying only on co-occurrence statistics, fail to capture hierarchical relationships, losing dependencies like disease co-occurrences and drug contraindications. (3) Redundancy across coding systems – Identical clinical concepts often appear under different codes across terminologies (e.g., ICD vs. SNOMED). Standard tokenization treats them as separate tokens, creating redundancy and complicating cross-system data integration. (4) Inefficiency in token storage – Expanding vocabulary sizes to accommodate medical codes results in bloated embedding tables that degrade computational efficiency, particularly for low-resource codes that appear infrequently but still require dedicated tokens. (5) Sparse and inconsistent usage – Many medical codes are rarely used or inconsistently documented, making it difficult for standard tokenizers to learn meaningful representations. Low-frequency codes suffer from poor embeddings, reducing performance on underrepresented conditions. (6) Lack of multimodal representations – Existing methods (Jiang et al., 2023b; Zhu et al., 2024; Xu et al., 2024) treat medical codes as isolated textual tokens, discarding graph-based relationships that encode essential links between diagnoses, treatments, and medications. A robust tokenizer must integrate both textual and relational information to fully represent medical codes.
Several models attempt to enrich the representations of medical codes by incorporating external knowledge from LLM (Jiang et al., 2023b; Zhu et al., 2024; Xu et al., 2024). Methods like GraphCare and RAW-EHR prompt LLMs to generate structured knowledge triplets of medical codes. Although effective in specific tasks, these approaches suffer from limited generalizability, a heavy reliance on knowledge generated by LLM, and a lack of a unified framework for handling various medical coding systems. Despite advances in medical representation learning, a unified tokenizer that integrates textual and structured relational knowledge across coding systems remains an open challenge.
Present work. We introduce MedTok, a multimodal medical code tokenizer that integrates textual descriptions and graph-based dependencies from biomedical ontologies (Figure 1). Unlike standard tokenization methods that treat medical codes as isolated textual tokens, MedTok captures both semantic meaning and structured relationships by encoding multiple modalities into a unified token space. MedTok operates in three stages. Multimodal encoding first extracts text embeddings from medical code descriptions and graph-based representations from biomedical knowledge graphs using separate encoders. Next, vector quantization maps both modalities into a shared token space, generating distinct text-informed and graph-informed token embeddings while preserving cross-modality interactions. Finally, optimization for expressivity ensures that token representations capture hierarchical relationships, semantic equivalence across different coding systems, and dependencies such as comorbidities and drug interactions.
We integrate MedTok into five EHR models and evaluate it in clinical and operational tasks that span the inpatient (MIMIC-III, MIMIC-IV) and outpatient (EHRShot) settings. These tasks include disease prediction, operational outcome modeling, drug recommendation, patient risk stratification, and operational outcomes. Key contributions:
  • Multimodal tokenization of medical codes – MedTok tokenizer jointly encodes both textual descriptions and graph-based representations of medical codes, enabling richer and structured embeddings.
  • Improved cross-system generalization – By incorporating ontological knowledge, MedTok bridges semantic gaps between different coding systems.
  • Demonstrated performance gains – Replacing standard EHR tokenizers with MedTok improves AUPRC by 4.10% on MIMIC-III, 4.78% on MIMIC-IV, and 11.30% on EHRShot, with the largest gains in drug recommendation tasks. MedTok is a general purpose tokenizer that can be integrated into any transformer-based model or system that requires tokenization. Beyond EHR models, we demonstrate its applicability in medical question-answering systems, further highlighting the benefit of optimized tokenization of structured medical data.
2 Related work
Domain-specific Tokenizers. Tokenizers tailored for specific domains have been employed to process various types of data, including language (Sennrich et al., 2016; Kudo & Richardson, 2018; Song et al., 2021; Wang et al., 2024b; Minixhofer et al., 2024), images (Zhou et al., 2022; Yu et al., 2022, 2024a; Zha et al., 2024), videos (Choudhury et al., 2024), graphs (Perozzi et al., 2024; Yang et al., 2024a), and molecular and material sciences (Fu et al., 2024; Tahmid et al., 2024; Qiao et al., 2024). While these tokenizers perform well within their respective domains, they are not directly applicable to medical codes, which contains specialized medical semantics. Medical codes reside in relation contexts and are accompanied by textual descriptions. Directly using the tokenizers for languages risks flattening the relationships among codes and failing to preserve the biomedical information. This will lead to fragmented tokenization of medical codes, resulting in loss of contextual information during encoding. Meanwhile, visual tokenizer typically focus on local pixel-level relationships, which are insufficient for capturing the complex semantics inherent in medical codes. Graph tokenizers are designed to encode structured information from graphs into a discrete token, then enabling LLMs to process relational and topological knowledge effectively. However, graph tokenizers may suffer from information loss when applied to graphs in other domains, making them less flexible and efficient for large, dynamic, and cross-domain graphs. In contrast, our MedTok tokenizer explicitly incorporates the relevant medical semantics by integrating textual descriptions with graph-based relational contexts.
Vector-Quantized Tokenizers. Tokenization strategies often vary according to the problem domain and data modality where recent work has highlighted the benefits of discrete tokenization (Du et al., 2024). This process involves partitioning the input according to a finite set of tokens, often held in a codebook (this concept is independent of medical coding despite the similar name), and the quantization process involves learning a mapping from input data to the optimal set of tokens according to a pre-defined objective such as reconstruction loss (Van Den Oord et al., 2017).
Recent work has highlighted the ability of vector quantized (VQ-based) tokenization to effectively compress semantic information(Gu et al., 2024). This approach is particularly successful for tokenizing inputs with an inherent semantic structure such as graphs (Yang et al., 2023a; Wang et al., 2024c), speech (Zeghidour et al., 2021; Baevski et al., 2019), and time (Yu et al., 2021) as well as complex tasks like recommendation retrieval (Wang et al., 2024d; Rajput et al., 2023; Sun et al., 2024) and image synthesis (Zhang et al., 2023; Yu et al., 2021).
Another significant advantage to VQ-based tokenization is the natural integration of multiple modalities. By learning a shared latent space across modalities, each modality can jointly modeled using a common token vocabulary (Agarwal et al., 2025; Yu et al., 2023). TokenFlow leverages a dual-codebook design that allows for correlations across modalities through a dual encoder (Qu et al., 2024).
Structured EHR, transformer-based, and foundation models. Structured EHR models leverage patient records to learn representations for clinical prediction and operational healthcare tasks. These models differ from medical LLMs (Singhal et al., 2025; Tu et al., 2024; Singhal et al., 2023), which are typically trained on free-text clinical notes (Jiang et al., 2023a) and biomedical literature rather than structured EHR data. BEHRT (Li et al., 2020) applies deep bidirectional learning to predict future medical events, encoding disease codes, age, and visit sequences using self-attention. TransformEHR (Yang et al., 2023b) adopts an encoder-decoder transformer with visit-level masking to pretrain on EHRs, enabling multi-task prediction. GT-BEHRT (Poulain & Beheshti, 2024) models intra-visit dependencies as a graph, using a graph transformer to learn visit representations before processing patient-level sequences with a transformer encoder. Other models enhance EHR representations with external knowledge. GraphCare (Jiang et al., 2023b) integrates large language models and biomedical knowledge graphs to construct patient-specific graphs processed via a Bi-attention Augmented Graph Neural Network. Mult-EHR (Chan et al., 2024) introduces multi-task heterogeneous graph learning with causal denoising to address data heterogeneity and confounding effects. ETHOS (Renc et al., 2024) tokenizes patient health timelines for transformer-based pretraining, achieving zero-shot performance. While these models focus on learning patient representations, MedTok serves a different role as a medical code tokenizer. It can be integrated into any structured EHR, transformer-based, or other foundation model, improving how medical codes are tokenized before being processed. Unlike these models, which rely on predefined tokenization schemes, MedTok optimizes the tokenization process itself.
3 Approach
Figure 2: MedTok is a general multimodal tokenizer of medical codes that can be integrated into any transformer-based model or a system that requires tokenization. ‘X-attn’ denotes a cross-attention module.
MedTok is a multimodal medical tokenizer that leverages both text descriptions and relational contexts of medical codes. MedTok operates as a tokenization function $f(\cdot)$ that maps a medical code $m\in\mathcal{M}$ to a sequence of elements $\mathcal{T}$ in the vocabulary $\mathcal{V}$ with a size of $N$ by leveraging both its textual definition $\mathcal{D}(m)$ and a subgraph $\mathcal{G}(m)$ extracted from a biomedical knowledge graph $G$. Here, $\mathcal{M}$ is a set of 617,490 medical codes from eight medical coding systems: ICD-9, ICD-10-CM, ICD-10-PCS, SNOMED CT, ATC, NDC, CPT, and RxNORM.
Problem definition. We formulate our problem as follows. Our goal is to train a multimodal tokenizer $f(\cdot)$ so that $\mathcal{T}=f(\mathcal{D}(m),\mathcal{G}(m))$, where $\mathcal{T}=[t_{1},t_{2},...,t_{T}]$ and $t_{i}\in\mathcal{V},1\leq i\leq T$. Then the generated $\mathcal{T}$ for medical code $m$ could be integrated to any EHR-based models $h(\cdot)$ and LMs models $p(\cdot)$ to perform predictive or generative tasks.
Overview. Fig. 2 illustrates the architecture of MedTok, which takes both the medical code description and contextual knowledge from biomedical KGs as input. MedTok takes two steps, multimodal tokenization and token packing.
3.1 Multimodal Tokenization
Given a medical code $m$, paired with its description $t$ and its biological subgraph $G$, MedTok first adopts the text encoder, denoted as $\mathrm{E}_{t}$ and the graph encoder, denoted as $\mathrm{E}_{g}$, to generate two embeddings: the text semantic embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{t}}$ for $t$ and the graph-level embedding $\mathbf{x}_{g}\in\mathbb{R}^{d_{g}}$ for $G$. These embeddings are computed as $\mathbf{x}_{t}=\mathrm{E}_{t}(t)$ and $\mathbf{x}_{g}=\mathrm{E}_{g}(G)$ for $G$, where $\mathrm{E}_{t}$ and $\mathrm{E}_{g}$ represent the text and graph encoders, respectively.
Modality-specific Embeddings. MedTok then adopts two linear projectors: $f_{t}:\mathbb{R}^{d_{t}}\rightarrow\mathbb{R}^{d}$ and $f_{g}:\mathbb{R}^{d_{g}}\rightarrow\mathbb{R}^{d}$, to generate modality-specific embeddings $\mathbf{e}_{t}^{s}\in\mathbb{R}^{d}$ and $\mathbf{e}_{g}^{s}\in\mathbb{R}^{d}$, respectively, where $\mathbf{e}_{t}^{s}=f_{t}(\mathbf{x}_{t})$, $\mathbf{e}_{g}^{s}=f_{g}(\mathbf{x}_{g})$, and $d$ is the dimension of specific embeddings.
Cross-modality Embeddings. Moreover, MedTok incorporates a cross-attention module to derive cross-modality embeddings $\mathbf{e}_{t}^{c}\in\mathbb{R}^{d}$ and $\mathbf{e}_{g}^{c}\in\mathbb{R}^{d}$, Specifically, the embedding $\mathbf{e}_{t}^{c}$ is computed as:
$\mathbf{e}_{t}^{c}=\mathrm{softmax}\left(\frac{\mathbf{W}_{q}\mathbf{x}_{t}(\mathbf{W}_{k}\mathbf{x}_{g})^{T}}{\sqrt{d}}\right)(\mathbf{W}_{v}\mathbf{x}_{g})$ (1)
Similarly, the embedding $\mathbf{e}_{g}^{c}$ is given by:
$\mathbf{e}_{g}^{c}=\mathrm{softmax}\left(\frac{\mathbf{W}_{q}\mathbf{x}_{g}(\mathbf{W}_{k}\mathbf{x}_{t})^{T}}{\sqrt{d}}\right)(\mathbf{W}_{v}\mathbf{x}_{t})$ (2)
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ represents the query, key, and value weight matrix.
Tokenization. After generating modality-specific and cross-modality embeddings, for each embedding, MedTok quantizes the embedding into $K$ tokens by querying a unified codebook $\mathbf{C}\in\mathbb{R}^{N\times d}$. The $K$ tokens are identified by the top $K$ nearest vectors in the codebook.
In detail, for any modality-specific or cross-modality embedding $\mathbf{e}_{:}$, its quantized tokens $\mathcal{I}(\mathbf{e}_{:})$ is formulated by:
$\mathcal{I}(\mathbf{e}_{:})=\mathrm{argmin}_{K}\left(\left\{\text{dist}(\mathbf{e}_{:},\mathbf{C}_{i})\right\}_{i=1}^{N}\right)$ (3)
where $\text{dist}(:,:)$ denotes the Euclidean distance, $|\mathcal{I}(\mathbf{e}_{:})|=K$, and $\mathbf{C}_{i}=\mathbf{C}[i,:]$. Then MedTok assigns a weight to each token $k\in\mathcal{I}(\mathbf{e}_{:})$ based on the distance between $\mathbf{e}_{:}$ and its corresponding vector $\mathbf{C}_{k}=\mathbf{C}[k,:]$. These weighted tokens are then summed together to obtain the quantized vector for $\mathbf{e}_{:}$, denoted as $\mathbf{\hat{e}}_{:}$, which is given by:
$\mathbf{\hat{e}_{:}}=\sum_{k\in\mathcal{I}(\mathbf{e}_{:})}-\mathrm{Softmax}(\mathrm{dist}(\mathbf{e}_{:},\mathbf{C}_{k}))*\mathbf{C}_{k}$ (4)
Following vector quantization conventions, we employ a straight-through gradient estimator: $\mathbf{e}_{:}=\mathrm{sg}[\mathbf{e}_{:}-\mathbf{\hat{e}}_{:}]+\mathbf{\hat{e}}_{:}$ where $\mathbf{sg}[\cdot]$ denotes the stop-gradient operation. The codebook learning objective is $\mathcal{L}(\mathbf{e}_{:},\mathbf{\hat{e}}_{:})=\|\mathrm{sg}[\mathbf{\hat{e}}_{:}]-\mathbf{e}_{:}\|_{2}^{2}+\alpha*\|\mathbf{\hat{e}}_{:}-\mathrm{sg}[\mathbf{e}_{:}]\|_{2}^{2}$, where $\alpha$ is a hyperparameter.
To preserve the distinctiveness of modality-specific and cross-modality embeddings, MedTok divides the entire codebook into three regions: a text-specific region, a graph-specific region, and a shared region, and then queries distinct regions of the codebook to generate their respective tokens and quantized vectors, which are represented by: ($\mathcal{I}(\mathbf{e}_{t}^{s})$, $\mathcal{I}(\mathbf{e}_{g}^{s})$, $\mathcal{I}(\mathbf{e}_{t}^{c})$, $\mathcal{I}(\mathbf{e}_{g}^{c})$) and ($\mathbf{\hat{e}}_{t}^{s}$, $\mathbf{\hat{e}}_{g}^{s}$, $\mathbf{\hat{e}}_{t}^{c}$, $\mathbf{\hat{e}}_{g}^{c}$). Consequently, the codebook learning objective is rewritten by:
$\mathcal{L}_{C}=\mathcal{L}(\mathbf{e}_{t}^{s},\mathbf{\hat{e}}_{t}^{s})+\mathcal{L}(\mathbf{e}_{g}^{s},\mathbf{\hat{e}}_{g}^{s})+\mathcal{L}(\mathbf{e}_{t}^{c},\mathbf{\hat{e}}_{t}^{c})+\mathcal{L}(\mathbf{e}_{g}^{c},\mathbf{\hat{e}}_{g}^{c})$ (5)
3.2 Token Packing
Unlike image-text paired data, where modalities have significant overlap, the two modalities (text and graph) of medical codes used in this work are more distinct yet complementary: the text focuses on clinical definitions, while the graph encodes domain-specific relationships not fully conveyed through text alone. Therefore, in addition to capturing shared information, it is crucial to extract modality-specific information during the tokenization process to ensure that the resulting tokens are highly informative.
Inspired by (Wang et al., 2024a), we optimize the obtained tokens ($\mathcal{I}(\mathbf{e}_{t}^{s})$, $\mathcal{I}(\mathbf{e}_{g}^{s})$, $\mathcal{I}(\mathbf{e}_{t}^{c})$, $\mathcal{I}(\mathbf{e}_{g}^{c})$) by separately optimizing the shared and modality-specific information across these tokens and their corresponding quantized vectors.
For shared information, given the tokens $\mathcal{I}(\mathbf{e}_{t}^{c})$ and $\mathcal{I}(\mathbf{e}_{g}^{c})$, the objective first is to adopt Kullback–Leibler (KL) divergence optimize them by ensuring their distance matrices $\mathrm{dist}(\mathbf{e}_{t}^{c},\mathbf{C})$ and $\mathrm{dist}(\mathbf{e}_{g}^{c},\mathbf{C})$ follow a similar distribution, as following:
$\mathcal{L}_{\text{KL}}=D_{\text{KL}}(\text{softmax}(\mbox{-}\mathrm{dist}(\mathbf{e}_{t}^{c},\mathbf{C}))\parallel\text{softmax}(\mbox{-}\mathrm{dist}(\mathbf{e}_{g}^{c},\mathbf{C})))$
Then optimize the quantized vectors $\mathbf{\hat{e}}_{t}^{c}$, $\mathbf{\hat{e}}_{g}^{c}$ to be highly informative about the other modality, while minimizing redundancy, as follows:
$\mathbf{\hat{e}}_{t}^{c*}=\arg\max_{\mathbf{\hat{e}}_{t}^{c}}\left(I(\mathbf{\hat{e}}_{t}^{c};\mathbf{e}_{g}^{c})-\beta\cdot I(\mathbf{\hat{e}}_{t}^{c};\mathbf{e}_{t}^{c}|\mathbf{e}_{g}^{c})\right)$ (6)
$\mathbf{\hat{e}}_{g}^{c*}=\arg\max_{\mathbf{\hat{e}}_{g}^{c}}\left(I(\mathbf{\hat{e}}_{g}^{c};\mathbf{e}_{t}^{c})-\beta\cdot I(\mathbf{\hat{e}}_{g}^{c};\mathbf{e}_{g}^{c}|\mathbf{e}_{t}^{c})\right)$ (7)
For specific information, given the tokens $\mathcal{I}(\mathbf{e}_{t}^{s})$ and $\mathcal{I}(\mathbf{e}_{g}^{s})$, the objective is to optimize these tokens by ensuring that the quantized vectors $\mathbf{\hat{e}}_{t}^{s}$, $\mathbf{\hat{e}}_{g}^{s}$ retain as much modality-specific information as possible, with minimal shared information, as follows:
$\mathbf{\hat{e}}_{t}^{s*}=\arg\max_{\mathbf{\hat{e}}_{t}^{s}}\left(I(\mathbf{\hat{e}}_{t}^{s},\mathbf{\hat{e}}_{g}^{c};\mathbf{e}_{t}^{s})-\lambda\cdot I(\mathbf{e}_{t}^{s};\mathbf{\hat{e}}_{g}^{c*})\right)$ (8)
$\mathbf{\hat{e}}_{g}^{s*}=\arg\max_{\mathbf{\hat{e}}_{g}^{s}}\left(I(\mathbf{\hat{e}}_{g}^{s},\mathbf{\hat{e}}_{t}^{c};\mathbf{e}_{g}^{s})-\lambda\cdot I(\mathbf{e}_{g}^{s};\mathbf{\hat{e}}_{g}^{c*})\right)$ (9)
Based on the derivation of (Wang et al., 2024a), the common tokens could be optimized by the combination of InfoNCE loss between $\mathbf{\hat{e}}_{t}^{c}$ and $\mathbf{\hat{e}}_{g}^{c}$, and the alignment loss between $\mathbf{e}_{t}^{c}$ and $\mathbf{e}_{g}^{c}$. Then the loss for packing shared information across two modalities is formulated by: $\mathcal{L}_{token}^{c}=\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{t}^{c},\mathbf{\hat{e}}_{g}^{c})+\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{g}^{c},\mathbf{\hat{e}}_{t}^{c})-2\beta\mathbb{E}_{\mathbf{e}_{t}^{c},\mathbf{e}_{g}^{c}}(\mathbf{e}_{t}^{c}\cdot\mathbf{e}_{g}^{c})$. Additionally, the modality-specific tokens could be optimized by the combination of InfoNCE loss between $\mathbf{\hat{e}}_{t}^{c}$, $\mathbf{\hat{e}}_{g}^{c}$ and the quantized vectors of their augmented data, and the orthogonal loss between $\mathbf{\hat{e}}_{t}^{c}$, $\mathbf{\hat{e}}_{g}^{c}$, and $\mathbf{e}_{t}^{c}$, $\mathbf{e}_{g}^{c}$, respectively. Then the loss for packing specific information across two modalities is formulated by: $\mathcal{L}_{token}^{s}=\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{t}^{c},\mathbf{\tilde{e}}_{t}^{c})+\lambda\mathcal{L}_{\mathrm{orthogonal}}(\mathbf{\hat{e}}_{t}^{c},\mathbf{e}_{t}^{c})+\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{g}^{c},\mathbf{\tilde{e}}_{g}^{c})+\lambda\mathcal{L}_{\mathrm{orthogonal}}(\mathbf{\hat{e}}_{g}^{c},\mathbf{e}_{g}^{c})$.
Finally, we combine the shared and specific losses to form the overall token packing loss, $\mathcal{L}_{token}=\mathcal{L}_{token}^{c}+\mathcal{L}_{token}^{s}$, to optimize our obtained tokens, where $\beta$ and $\lambda$ are hyperparameters set to be equal. This approach enables MedTok to leverage both modality-shared and modality-specific information.
3.3 Training and Inference for MedTok
Training Stage. During the training stage, MedTok is trained by the sum of codebook loss $\mathcal{L}_{C}$, KL divergency loss $\mathcal{L}_{KL}$, loss for packing common and specific tokens across two modalities $\mathcal{L}_{token}$, where $\mathcal{L}=\mathcal{L}_{C}+\mathcal{L}_{KL}+\mathcal{L}_{token}$.
Inference Stage. After pre-training, MedTok can be seamlessly integrated into any model or pipeline dealing with medical codes, providing unified medical tokens for downstream tasks.
4 Experiments
Medical coding systems. We collected a total of 617,490 medical codes from eight commonly used coding systems: ICD-9 (Organization et al., 1988), ICD-10-CM (Fung et al., 2020), ICD-10-PCS (Averill et al., 2001), SNOMED CT (Donnelly et al., 2006), ATC (Miller & Britt, 1995), NDC (Palmer, 2006), CPT (Dotson, 2013), and RxNORM (Nelson et al., 2011), as shown in Table 1. These codes cover various events, including procedures, diagnoses, and medications. Each code is paired with a textual description from official documents and a subgraph from PrimeKG (Chandak et al., 2023). Details for data pre-processing are available in Appendix A.
Code System Count Code System Count
SNOMED 303,325 ICD9 18,365
ICD10-CM 81,184 CPT 10,602
RxNorm 81,151 ATC 6,659
ICD10-PCS 61,644 NDC 54,560
Table 1: Summary of the dataset’s code systems distribution.
Patient EHR datasets. We used three publicly available EHR datasets: MIMIC-III (Johnson et al., 2016), MIMIC-IV (Johnson et al., 2024), and EHRShot (Wornow et al., 2023). MIMIC-III and MIMIC-IV are in-patient datasets with medical records for ICU patients, while EHRShot is a dataset containing longitudinal medical records that include both out-patients and ICU/ED patients. MIMIC datasets include NDC medications and ICD-9 / ICD-10 codes for diagnoses and procedures. In contrast, EHRShot mainly uses RxNorm codes for medications, SNOMED codes for diagnoses, and CPT, SNOMED, ICD-9, and ICD-10 codes for procedures. Table 2 summarizes the statistics of three EHR datasets.
#patients #visits #visits/patient #events/patient
MIMIC-III 35,707 44,399 1.24 51.14
MIMIC-IV 123,488 232,263 1.88 70.33
EHRShot 6,739 921,499 136.74 6182.17
Table 2: Statistics of EHR datasets.
Model Task 1: MT+ Task 2: RA($<$15 days)+ Task 3: LOS Task 4: Pheno Task 5: DrugRec
MIMIC-III MIMIC-IV MIMIC-III MIMIC-IV MIMIC-III MIMIC-IV MIMIC-III MIMIC-IV MIMIC-III MIMIC-IV
AUPRC AUPRC AUPRC AUPRC AUPRC AUPRC AUPRC AUPRC AUPRC AUPRC
ETHOS 0.617 (0.010) 0.282 (0.001) 0.421 (0.007) 0.648 (0.005) N/A N/A N/A N/A 0.104 (0.008) 0.131 (0.005)
+ MedTok 0.634 (0.020) 0.412 (0.030) 0.463 (0.017) 0.690 (0.007) N/A N/A N/A N/A 0.170 (0.014) 0.240 (0.012)
GT-BEHRT 0.160 (0.037) 0.028 (0.004) 0.612 (0.058) 0.586 (0.070) 0.230 (0.010) 0.103 (0.001) 0.423 (0.002) 0.493 (0.005) 0.715 (0.002) 0.736 (0.007)
+ MedTok 0.193 (0.046) 0.034 (0.005) 0.623 (0.052) 0.609 (0.064) 0.287 (0.039) 0.114 (0.003) 0.459 (0.028) 0.512 (0.006) 0.740 (0.004) 0.783 (0.010)
MulT-EHR 0.136 (0.021) 0.120 (0.003) 0.574 (0.008) 0.515 (0.007) 0.176 (0.018) 0.118 (0.032) 0.460 (0.012) 0.498 (0.001) 0.523 (0.008) 0.445 (0.027)
+ MedTok 0.156 (0.025) 0.141 (0.013) 0.585 (0.016) 0.565 (0.002) 0.198 (0.011) 0.136 (0.030) 0.480 (0.002) 0.504 (0.001) 0.571 (0.006) 0.465 (0.003)
TransformEHR 0.207 (0.012) 0.042 (0.012) 0.527 (0.030) 0.518 (0.012) 0.132 (0.021) 0.119 (0.001) 0.469 (0.022) 0.507 (0.007) 0.533 (0.030) 0.612 (0.046)
+ MedTok 0.246 (0.044) 0.058 (0.007) 0.568 (0.036) 0.525 (0.017) 0.159 (0.031) 0.121 (0.002) 0.513 (0.024) 0.518 (0.012) 0.580 (0.035) 0.661 (0.092)
BEHRT 0.163 (0.037) 0.028 (0.003) 0.529 (0.053) 0.514 (0.015) 0.232 (0.015) 0.112 (0.003) 0.587 (0.004) 0.493 (0.006) 0.539 (0.013) 0.778 (0.014)
+ MedTok 0.220 (0.025) 0.032 (0.006) 0.574 (0.040) 0.515 (0.005) 0.251 (0.030) 0.137 (0.004) 0.603 (0.008) 0.504 (0.006) 0.558 (0.006) 0.792 (0.007)
Improvement (%) +3.32% 3.54% 3.00% 2.46% 3.13% 1.40% 2.90% 1.18% 4.10% 4.78%
  • $+$: imbalanced binary classification; *: multi-class classification, macro-averaged; $\circ$: multi-label classification; N/A indicates that the model was not configured for this task.
Table 3: The results of MedTok with all baseline models across five tasks on two in-patient datasets.
Baselines. To evaluate MedTok for tasks related to medical codes, we consider two types of tokenizers and five models based on EHR as baselines. The first type of tokenizer is text-based (e.g., bert-base-uncased (Devlin et al., 2018)), while the second is graph-based (e.g., VQGraph (Yang et al., 2024b)). Five EHR-based models are ETHOS (Renc et al., 2024), GT-BEHRT (Poulain & Beheshti, 2024), MulT-EHR (Chan et al., 2024), TransformEHR (Yang et al., 2023b), and BEHRT (Li et al., 2020). Details on implementation can be found in the Appendix B.
Evaluation setup. We consider two evaluation setups:
  • In-patient evaluation: This setting combines the MedTok tokenizer with patient prediction models, using two in-patient datasets that include individuals admitted to a hospital. The evaluation encompasses five tasks: 1 mortality prediction (MT), 2 readmission prediction (RA), 3 length-of-stay prediction (LOS), 4 phenotype prediction (Pheno), and 5 drug recommendation (DrugRec). The first three tasks focus on predicting a patient’s future health status using their historical medical records. Phenotype prediction involves the identification of the phenotype of a patient’s disease based on their medical history. We identified 24 phenotypes for diseases in MIMIC-III and MIMIC-IV, as follows (Harutyunyan et al., 2019). Drug recommendation aims to suggest appropriate medications for a patient, considering their historical medical records and the diseases identified during their current visit. For drug recommendation, we focus on five specific drug candidates, including Vancomycin, Levofloxacin, Heparin Sodium, Metoprolol, and Atorvastatin, rather than considering the entire range of available medications. AUPRC is adopted to evaluate the model’s performance on the above classification tasks.
  • Out-patient evaluation: We evaluate MedTok together with patient prediction models on a dataset of patients who are not admitted to a hospital and consider two categories of tasks: 1 Operational Outcomes (OO), and 2 new diagnosis assignments (ND), following (Wornow et al., 2023). The OO includes MT, RA, and prolonged LOS. The new diagnosis assignments are used to predict the first diagnosis of a disease. Details of task definitions are in Appendix C.
4.1 MedTok tokenizer with in-patient EHR models
Table 3 presents the AUPRC values for each baseline and their integration with our MedTok for five tasks in two in-patient datasets. Compared to baselines that treat each medical code as an individual token, integrating our MedTok consistently improves performance across all five tasks, achieving an average improvement of 3.29% on MIMIC-III and 2.67% on MIMIC-IV. This improvement comes from more informative tokens generated by MedTok, which strengthen the EHR-based models. Among five tasks, MedTok demonstrates the most significant impact on drug recommendation tasks, highlighting the value of incorporating prior knowledge into our tokenizer.
To further assess the effectiveness of MedTok, we compare it against two tokenization methods: the text-based BERT tokenizer and the graph-based VQGraph tokenizer. Figure 3 presents the performance of each tokenizer when integrated with a Transformer-based EHR model (TransformEHR) across five tasks on two in-patient datasets. MedTok consistently outperforms BERT and VQGraph in all tasks and datasets, demonstrating the superiority of its tokenization strategy.
Model Task 1: Operational Outcomes (OO) Task 2: Assignment of New Diagnoses (ND)
Long LOS RA (<{<}<15 days) MT Hypertension Hyperlipidemia Pancreatic Cancer Acute MI
AUPRC AUPRC AUPRC AUPRC AUPRC AUPRC AUPRC
ETHOS NA 0.079 (0.017) 0.102 (0.018) 0.166 (0.020) 0.155 (0.031) 0.056 (0.006) 0.093 (0.011)
+ MedTok NA 0.128 (0.025) 0.339 (0.010) 0.175 (0.019) 0.163 (0.025) 0.056 (0.013) 0.104 (0.017)
GT-BEHRT 0.714 (0.021) 0.115 (0.012) 0.239 (0.012) 0.303 (0.018) 0.239 (0.007) 0.044 (0.008) 0.015 (0.008)
+ MedTok 0.739 (0.025) 0.154 (0.013) 0.444 (0.015) 0.360 (0.012) 0.441 (0.005) 0.074 (0.010) 0.031 (0.015)
MulT-EHR 0.539 (0.025) 0.125 (0.014) 0.397 (0.016) 0.218 (0.005) 0.243 (0.005) 0.022 (0.008) 0.017 (0.003)
+ MedTok 0.571 (0.015) 0.188 (0.021) 0.444 (0.012) 0.226 (0.006) 0.254 (0.021) 0.037 (0.015) 0.028 (0.014)
TransformEHR 0.652 (0.023) 0.197 (0.016) 0.344 (0.030) 0.376 (0.018) 0.305 (0.021) 0.053 (0.006) 0.025 (0.006)
+ MedTok 0.675 (0.018) 0.243 (0.016) 0.379 (0.034) 0.413 (0.026) 0.333 (0.018) 0.082 (0.012) 0.052 (0.017)
BEHRT 0.582 (0.032) 0.332 (0.022) 0.389 (0.018) 0.233 (0.027) 0.251 (0.019) 0.036 (0.008) 0.013 (0.031)
+ MedTok 0.723 (0.028) 0.397 (0.036) 0.431 (0.017) 0.287 (0.018) 0.302 (0.015) 0.057 (0.012) 0.036 (0.015)
Improvement (%) +5.52% +5.24% +11.32% +3.30% +6.00% +1.90% +1.76%
Table 4: The results of MedTok with all baseline models across two tasks on the EHRShot dataset.
Figure 3: The AUPRC values of three types of tokenizers on in-patient and out-patient datasets, where OO means Operational Outcomes and ND means assignment of new diagnoses.
4.2 MedTok tokenizer with out-patient EHR models
Table 4 presents the AUPRC values for each baseline and its integration with MedTok across two task types on the out-patient HRShot dataset. The results reveal that our tokenizer has the most significant impact on mortality prediction in Operational Outcomes, achieving an average improvement of 11.30%. It also significantly improves the detection of new diagnoses of Acute MI, with an average improvement of 8.80%. As shown in Fig. 3, a comparison of three types of tokenizers further demonstrates the effectiveness of MedTok in integrating both graph and textual modalities. Additionally, when comparing performance across two in-patient datasets, we observe that MedTok is particularly beneficial for longitudinal data.
4.3 Ablation studies
In the ablation studies, to eliminate potential bias from different model architectures, we integrate MedTok with a vanilla Transformer-based model (e.g., TransformEHR) to examine the impact of the adopted modalities and the vocabulary size in MedTok on performance.
Figure 4: The AUPRC values obtained by removing the text and graph modalities across all tasks on two in-patient datasets and one out-patient dataset.
Effects of modalities on MedTok. To evaluate the impact of the two modalities (text, graph) used in MedTok —medical code definitions and biological subgraphs derived from a biomedical knowledge graph—we assess its performance by removing the text and graph modalities separately. As shown in Fig. 4, MedTok, when leveraging both modalities, achieves the best performance across all tasks on three datasets. By comparing the performance of MedTok without the graph modality and MedTok without the text modality, we observe that both modalities contribute significantly to EHR-based prediction tasks. The graph modality benefits drug recommendation and new disease detection tasks, while the text modality proves essential for readmission prediction on MIMIC-III and operational outcomes in EHRShot. These findings emphasize the importance of incorporating the underlying information linked to medical codes.
Effect of codebook size $N$. We further evaluate the impact of the codebook size on the performance of MedTok by training it with varying sizes and assessing its effectiveness across three distinct datasets integrated with TransformEHR. Fig. 5 presents the results for various codebook sizes across all tasks on the three datasets. The performance trends observed on MIMIC-III and MIMIC-IV are quite consistent, demonstrating a clear pattern where increasing the codebook size enhances the model’s performance. Specifically, the highest average performance is achieved when the codebook size is set to $N=12,000$, indicating that this size strikes an optimal balance between sufficient coverage of the medical vocabulary and avoiding overfitting.
In contrast, when analyzing the performance on EHRShot, a dataset consisting of patients with longer visit histories than those in MIMIC-III and MIMIC-IV, we observe that MedTok benefits from a larger codebook size. For EHRShot, the highest average performance is achieved when the codebook size is increased to $N=21,000$. This suggests that for datasets with more extensive patient visit histories, a larger codebook may be more effective in capturing the underlying complexity of the medical information, thus improving the model’s predictive capabilities.
4.4 Using MedTok tokenizer for medical QA
MedTok demonstrates strong performance in EHR-based tasks, as shown in Tables 3-4. To further assess its capabilities, we explore its effectiveness in a generation task, specifically multiple-choice medical question answering (MedicalQA), where the goal is to select the correct answer to a given clinical question (Singhal et al., 2023). We evaluate whether MedTok enhances few-shot learning in MedicalQA by integrating its tokenized representations with a LLM (LLaMA3.1-8B (Dubey et al., 2024)). MedTok-generated tokens are used as prefix tokens, which provide structured medical context before the main input, allowing the LLM to incorporate additional domain knowledge.
For this evaluation, we use the MedDDx dataset (Su et al., 2024), which contains questions at three difficulty levels: Basic, Intermediate, and Expert. The process consists of three steps: (1) Disease code mapping – Extract disease mentions from each question and retrieve their corresponding medical codes. (2) Tokenization via MedTok – Convert medical codes into structured tokens using MedTok. (3) Prefix Token Fine-Tuning – Fine-tune LLaMA3.1-8B using MedTok tokens as prefix inputs before the question text. We fine-tune the model on 617 intermediate-level questions and evaluate performance on 227 expert-level and 158 basic-level questions. The results in Figure 5 show an accuracy improvement of 0.3% on MedDDx-Basic and 2.3% on MedDDx-Expert, suggesting that MedTok can enhance medical QA when used as a structured prefix representation.
Figure 5: A, The AUPRC values of MedTok with different codebook size $N$; B, The accuracy of LLaMA3.1-8B vs. MedTok +LLaMA3.1-8B on two MedDDx medical QA datasets.
5 Conclusion
Tokenizing medical codes is a critical yet challenging step in developing foundation models for electronic health records (EHRs). Existing tokenizers treat medical codes as isolated textual units, failing to capture their structured relationships within large-scale medical ontologies. With more than 600,000 codes that span multiple terminologies, standard tokenization methods struggle to scale while preserving the rich semantic and relational context necessary for downstream clinical and operational tasks. We introduced MedTok, a multimodal tokenizer of medical codes that integrates textual definitions and relational ontologies of medical codes to create a unified token representation. MedTok applies vector quantization to encode both modalities in a structured token space, preserving cross-modality relationships. We integrated MedTok with five EHR models, evaluating its impact across inpatient (MIMIC-III, MIMIC-IV) and outpatient (EHRShot) settings, as well as in fine-tuning a medical question-answering system. Our results establish MedTok as a generalizable tokenizer for medical codes, shedding light on how optimizing the tokenization process can benefit medical foundation models.
Impact Statement
This work presents a tokenizer for medical codes designed to assist other models in better encoding semantic knowledge. While our tokenizer complements other models, it does not directly raise ethical concerns. Instead, it enhances the trustworthiness of these models by providing text and graph-based information as references for tokenization. In future work, we will explore ways to better integrate our tokenizer with other models to further improve both their performance and trustworthiness.
Acknowledgement
We sincerely appreciate the valuable discussions with Intae Moon and Zhenglun Kong, whose insights and feedback have contributed to the development of this work. We gratefully acknowledge the support of NIH R01-HD108794, NSF CAREER 2339524, US DoD FA8702-15-D-0001, ARPA-H BDF program, awards from Chan Zuckerberg Initiative, Bill & Melinda Gates Foundation INV-079038, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, Roche Alliance with Distinguished Scientists, Sanofi iDEA-iTECH, Pfizer Research, John and Virginia Kaneb Fellowship at Harvard Medical School, Biswas Computational Biology Initiative in partnership with the Milken Institute, Harvard Medical School Dean’s Innovation Fund for the Use of Artificial Intelligence, Harvard Data Science Initiative, and Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders.
Conflict of Interest
Faryad Sahneh is a Sanofi employee and may hold shares and/or stock options in the company. Other authors declare no conflict of interest.
Appendix A Data preprocessing details
A.1 Medical Codes Dataset Creation
The medical codes dataset consists of medical codes, their descriptions, and associated knowledge subgraphs, encompassing eight commonly used health coding systems: ICD-9-CM (procedures and diagnoses), ICD-10-CM, ICD-10-PCS, NDC (National Drug Codes), SNOMED CT, ATC (Anatomical Therapeutic Chemical Classification), CPT (Current Procedural Terminology), and RxNorm. All code lists were obtained from official sources. Specifically, ICD-9 and ICD-10 (CM and PCS) were sourced from the CMS website; NDC codes from the U.S. Food and Drug Administration (FDA) database; and CPT (Level I HCPCS) from the Physician Fee Schedule (PFS) Relative Value Files at CMS. SNOMED CT, RxNorm (active codes only), and ATC were downloaded via the National Library of Medicine (NLM), part of the National Institutes of Health (NIH).
A.1.1 Medical Codes Knowledge Graphs Creation
In the final dataset, each medical code is linked to a knowledge graph capturing relevant medical insights and relationships. We constructed these subgraphs in two steps: mapping each code to one or more nodes in the PrimeKG knowledge graph; and extracting node-centered subgraphs to represent the code’s associated knowledge and connections. To facilitate mapping, we leveraged several external resources, notably the UMLS database and MONDO Disease Ontology files. Medical codes were first mapped to Concept Unique Identifiers (CUIs) in the UMLS database, then linked to PrimeKG nodes via a custom UMLS-to-PrimeKG file. Because PrimeKG includes MONDO annotations, we also aligned medical codes to MONDO terms using the mondo.owl file, thus achieving direct integration with PrimeKG nodes. Additionally, a custom entity linker was employed to enhance coverage by translating medical codes into descriptive text (via PyHealth’s MedCode InnerMap) and matching these descriptions to PrimeKG node names. When exact matches were unavailable, we resorted to an NLP-based linker (SciSpacy with UMLS) to measure semantic similarity. For drug codes, the rxnav.nlm.nih.gov API was used to map RxNorm codes to ATC identifiers, which were then associated with DrugBank entities through a predefined ATC-to-DrugBank mapping.
A.1.2 Medical Codes Textual Definition Creation
Initially, each medical code’s description was taken from its official source. For medication codes (e.g., NDC) where the original text was sparse, additional details were derived from attributes such as trade name, proprietary name, and pharmacological classification. These preliminary definitions were then refined and enriched using ChatGPT-4 (turbo), with prompts tailored to each coding system but sharing a common goal of elaborating on clinical uses (for drugs), procedural steps (for procedures), or mechanistic and clinical context (for diagnoses).
Appendix B Implementation details
B.1 Experimental environments
Hardware. MedTok is training on a machine equipped with 4 NVIDIA H100. All experiments were conducted with 1 NVIDIA H100.
Software.
We implement MedTok using Python 3.9.19, PyTorch 2.3.1, Transformers 4.43.1, and Tokenizers 0.19.1. All LMs and LLMs adopted in this study are downloaded from Hugging Face, except for OpenAI models.
B.2 Details in MedTok training
MedTok is trained on 4 NVIDIA H100 GPUs by using the loss defined in the Section 3.2. During the training stage, we set the training step as 3000 with a global batch size of 1024, the dimension of quantized vectors is 64. In terms of the models’ weights, we freeze the text encoder in MedTok and the graph encoder is trainable during the training stage.
B.3 Implementation details of baseline models
All results presented in this study were obtained using the same machine on which the MedTok was trained.
ETHOS experiments were conducted using the authors’ original repository. For each experimental setting, three models were trained on the MIMIC-IV dataset with different random seeds, and their predictions were averaged during inference to ensure robustness. In the ”MedTok + ETHOS” configuration, the original vocabulary was extended to incorporate MedTok’s tokens for diagnoses, procedures, and prescriptions. The lab measurements were excluded from the analysis. Training and dataset splitting on MIMIC-IV adhered to the methodology outlined in the ETHOS paper (Renc et al., 2024). During inference, the number of generated tokens was limited to 2048, and the timeline duration was adjusted based on the specific task: fifteen days for readmission, two weeks for mortality, and up to six months for other tasks. Each model was executed five times, and the resulting predictions were averaged to produce a continuous output, as described in the ETHOS study. Inference on the MIMIC-III dataset was performed on the entire dataset, excluding BMI, ICU stay tables, blood pressure, and lab data. For the EHRShot dataset, inference was conducted on the full dataset for mortality and disease-related tasks, and on randomly selected, stratified samples of ten thousand instances for other tasks.
As for the other baselines adopted in this work, we first downloaded their code and deploy these models on our working machine. For BEHRT and GT-BEHRT, we re-trained it in an end-to-end way and integrates the tokens for time, visit, and patient’s info as that in their original work. For MulT-EHR, we first pre-train it on MIMIC-III, MIMIC-IV, and EHRShot, respectively, to get the embedding of medical codes, and next fine-tune it on multi-task learning. For the MedTok +, we use our token embeddings to initialize the nodes or tokens the original work adopted and then train or pre-train the model. It should be noted that we adopt a unified epoch number for all baselines, which is 50.
Appendix C Task definitions and data preparation under in-patient setting
C.1 Mortality prediction
Task definition. Mortality (MT) prediction estimates the mortality label of the subsequent visit for each sample, with the last sample dropped. Formally,
$f:(v_{1},v_{2},\ldots,v_{t-1})\;\to\;y[v_{t}],$
where $y[v_{t}]\in\{0,1\}$ is a binary label indicating the patient’s survival status recorded in visit $v_{t}$.
C.2 Readmission prediction
Task definition. Readmission prediction checks if the patient will be readmitted to the hospital within $\sigma$ days. Formally, $f:(v_{1},v_{2},\ldots,v_{t-1})\;\to\;y\bigl{[}\tau(v_{t})-\tau(v_{t-1})\bigr{]}$, where $y\in\{0,1\}$ and $\tau(v_{t})$ denotes the encounter time of visit $v_{t}$. Specifically,
$y\bigl{[}\tau(v_{t})-\tau(v_{t-1})\bigr{]}\;=\;\begin{cases}1&\text{if }\tau(v_{t})-\tau(v_{t-1})\leq\sigma,\\ 0&\text{otherwise}.\end{cases}$
In our study, we set $\sigma=15$ days.
C.3 Length-of-Stay (LOS) prediction
Task definition. Length-of-Stay (LOS) prediction follows the formulation of Harutyunyan et al., estimating ICU stay length for each visit. Formally, $f:(v_{1},v_{2},\ldots,v_{t})\to y[v_{t}]$, where $y[v_{t}]\in\mathbb{R}^{1\times C}$ is a one-hot vector indicating its class among $C$ possible categories. We define 10 classes, $\{0,1,\ldots,7,8,9\}$, representing the following durations: 0 for one day or less, 1-7 for within one week, 8 for one to two weeks, and 9 for at least two weeks.
C.4 Phenotype prediction
Task definition. Phenotype prediction aims to classify which acute care conditions are present in a given patient record: $f:(v_{1},v_{2},...,v_{t})\rightarrow y[v_{t}]$, where $y[v_{t}]\in\mathbb{R}^{1\times C}$ is a one-hot vector indicating its class among $C$ possible categories. This task is a multilable classification problem with macro-averaged AUC-ROC being the main metric.
C.5 Drug recommendation
Task definition. Drug recommendation aims to recommend drugs for a patient according to the patient’s visit history and diagnosis in current visit: $f:(v_{1},v_{2},...,v_{t})\rightarrow y[v_{t}]$, where $y[v_{t}]\in\mathbb{R}^{1\times C}$ is a one-hot vector indicating its class among $C$ possible categories. This task is a multilable classification problem with macro-averaged AUC-ROC being the main metric.
Data preprocessing. In this study, we adopted a data preprocessing approach similar to that used in previous research (https://doi.org/10.1038/s41597-019-0103-9), which defined 25 acute care conditions. Each diagnosis code was mapped to one of these 25 phenotype categories. Since ICD-9 codes in MIMIC-III are associated with hospital visits rather than specific ICU stays, we linked diagnoses to ICU stays using the hospital admission identifier. To reduce ambiguity, we excluded hospital admissions involving multiple ICU stays, ensuring that each diagnosis corresponded to a single ICU stay per admission. It’s important to note that our phenotype classification was retrospective; we analyzed the complete ICU stay before predicting the presence of specific diseases. In this study, we adopted a data preprocessing approach similar to that used in previous research (https://doi.org/10.1038/s41597-019-0103-9), which defined 25 acute care conditions. Each diagnosis code was mapped to one of these 25 phenotype categories. Since ICD-9 codes in MIMIC-III are associated with hospital visits rather than specific ICU stays, we linked diagnoses to ICU stays using the hospital admission identifier. To reduce ambiguity, we excluded hospital admissions involving multiple ICU stays, ensuring that each diagnosis corresponded to a single ICU stay per admission. It’s important to note that our phenotype classification was retrospective; we analyzed the complete ICU stay before predicting the presence of specific diseases.
C.6 Out-patient Setting
Under this setting, we adopt two types of tasks in EHRShot, including operational outcomes prediction and assignment of new diagnosis. In the field of operational outcomes, we follow the same task definitions in long length of stay prediction, which only consider if a patient stay in the hospital less than 7 days or more than 7 days. In terms of readmission task, we set the time window as 15 days, which is the same as that under in-patient setting. We also add another operational outcome tash, which is mortality prediction. The definition of mortality prediction is the same as that under the in-patient setting.